# Predicting and Understanding Turn-Taking Behavior in Open-Ended Group Activities in Virtual Reality

### v1

@@ Abstract.
In networked virtual reality (VR), user behaviors, individual differences, and group dynamics can serve as important signals into future speech behaviors, such as who the next speaker will be and the timing of turn-taking behaviors. The ability to predict and understand these behaviors offers opportunities to provide adaptive and personalized assistance, for example helping users with varying sensory abilities navigate complex social scenes and instantiating virtual moderators with natural behaviors. In this work, we predict turn-taking behaviors using features extracted based on social dynamics literature. We discuss results from a large-scale VR classroom dataset consisting of 77 sessions and 1660 minutes of small-group social interactions collected over four weeks. In our evaluation, gradient boosting classifiers achieved the best performance, with accuracies of 0.71–0.78 AUC (area under the ROC curve) across three tasks concerning the “what”, “who”, and “when” of turn-taking behaviors. In interpreting these models, we found that group size, listener personality, speech-related behavior (e.g., time elapsed since the listener’s last speech event), group gaze (e.g., how much the group looks at the speaker), as well as the listener’s and previous speaker’s head pitch, head y-axis position, and left hand y-axis position more saliently influenced predictions. Results suggested that these features remain reliable indicators in novel social VR settings, as prediction performance is robust over time and with groups and activities not used in the training dataset. We discuss theoretical and practical implications of the work.
@@ 1. Introduction
Effective and organized social interactions require its social actors to accurately infer speaking intention and take turns accordingly. In most platforms for facilitating computer mediated-communication (CMC), users retain a high level of autonomy over when and how they speak, for example utilizing features such as the muting and unmuting button during videoconferencing. However, along with the benefits of user control comes the responsibility for individuals to correctly infer social group dynamics. Such inferences can be difficult to navigate as users utilize digital platforms such as video conferencing and social virtual reality (VR), where the level of immersion and representation of users (e.g., nonverbal behavior, avatar appearance) can differ from one another (Steinicke et al., 2020; Abdullah et al., 2021). In immersive social interactions, differences in avatar representation and characteristics of social VR platforms can yield adverse outcomes such as unwanted overlap in speech (Barreda-Ángeles et al., 2023; Williamson et al., 2021), lack of control for users with physical disabilities (Maloney et al., 2020), misperceived and absence of nonverbal cues (Maloney et al., 2020; Moustafa and Steed, 2018), and difficulty when gauging the end of speaking turns (Moustafa and Steed, 2018; Williamson et al., 2021). For users with visual and audio disabilities, it can also be challenging for them to engage in seamless interactions based solely off of perceptual and auditory feedback (Kushalnagar and Vogler, 2020; Anderson, 2021). Paradoxically, as immersive communication evolves to be hardware-agnostic, with users interacting across traditional screen devices and mixed-reality headsets (Qiu et al., 2023), the mismatch in social cues delivered and perceived by users across different devices can present further complications for users to navigate socially.
@@
Being able to predict user intentions for actions such as turn taking can alleviate these challenges through real-time interventions. Social VR platforms can implement personalized assistance for users who are struggling to interpret nonverbal cues under technological constraints (e.g., joining large-group VR meetings though audio feed) or are different in physical abilities (e.g., auditory and visual abilities). In immersive social settings requiring more structure such as classroom and focus groups, knowing who will likely speak next and what nonverbal cues precede speech can further improve social dynamics. Practitioners can instantiate virtual agents to moderate social VR scenes with natural nonverbal behaviors that do not cut off the current speaker and more broadly design mechanisms for smoothly guiding the conversations. Failure to accurately infer social dynamics (e.g., speaking intentions) can yield unwanted interruptions and lower the perceived ease of use of these intelligent systems (Maier et al., 2022). As such, we argue that it is important to (1) study whether we can robustly predict turn-taking behaviors in diverse VR settings and (2) how nonverbal and verbal behaviors as well as individual and group characteristics influence turn-taking predictions.
@@
In this work, we investigated turn-taking behaviors in VR and leveraged the medium’s fine-grain tracking data on motion and verbal behaviors. Using the VR motion data, researchers have implemented pipelines for automatically deriving features such as head movement (Miller et al., 2023a) and gaze (Abdullah et al., 2021)), both of which have been previously useful for building models for predicting speech behaviors in face-to-face social interactions (Ishii et al., 2017, 2016b). We further leveraged VR’s ability to place groups into varying virtual spaces irrespective of the users’ physical locations, which enabled us to model and predict turn-taking behaviors across different groups, social activities, and spatial contexts over time. Studying diverse social interactions at scale is important as building predictive models that are robust to unseen groups, activities, and time suggests higher practicality and reliability when deploying such models to novel social settings. Concretely, our work aims to answer three research questions, namely
@@
Research Question 1 (RQ1): Can we predict turn-taking behaviors in VR open-ended group activities from features extracted from individual- and group- level differences, speech-related behavior, and motion tracking data?
@@
Research Question 3 (RQ3): How are the extracted features related to turn-taking predictions and performance? Specifically, what features are strongly related to model performance, and how are they associated with predictions of turn-taking behavior?
To do this, we built models predicting speech dynamics, specifically focusing on three tasks that capture the “what”, “who”, and “when” of turn-taking behaviors. We formulated features based on tracking data and individual and group characteristics based on past literature on social dynamics. We evaluated performance across 77 sessions and 1660 minutes of open-ended VR group discussions spanning four weeks, each with three to four university students. Notably, we found that gradient boosting classifiers achieved the best accuracies, predicting turn transitions and identifying the new speaker with 0.75–0.78 AUC (i.e., area under the ROC curve), and differentiating between moments immediately preceding turn transitions and those sampled at prior moments with accuracies of 0.71–0.72 AUC. These results demonstrated that we were able to predict turn-taking behaviors with accuracies considerably higher than prediction by chance (i.e., 0.50 AUC) and highlighted the potential non-linear relationships and interactions between our extracted features.
Additional feature analyses on the models revealed the importance of listener personality, group size, speech-related behavior (e.g., preceding speaker sequence, time to listener’s most recent speaking turn), group gaze behavior, head pitch and y-axis position, and left hand y-axis position. Our models showed comparable results when evaluated on time, activities, and groups not seen during training, which demonstrated the reliability of these turn-taking indicators in novel social settings. From these results, we highlight their theoretical relevance and outline practical implications for how practitioners can leverage behavioral predictions to implement real-time interventions and adaptive assistance to facilitate effective VR social interactions.
In summary, we make the following contributions. First, we formulated features based on past literature on social dynamics and leveraged both the VR tracking data and individual and group characteristics. Using a large-scale VR dataset with 1660 minutes of open-ended activities collected over time, we then demonstrated the feasibility and robustness of using these features to predict VR turn-taking behaviors. Finally, through interpreting how individual features influence model performance and prediction probabilities, we present theoretical and practical insights that we hope will motivate future VR research on social dynamics, particularly given its potential to enable real-time intervention and adaptive assistance in immersive social settings.
@@ 2. Related Work
In this section, we describe literature related to ours, namely those related to nonverbal behaviors in social interactions, modeling behaviors and individual- and group- level differences, and social interactions across individuals and groups. For each topic, we first provide an overview of prior work and then contextualize our work and contributions within the area of study.
@@ 2.1. Nonverbal Behavior in Social Interactions
Nonverbal behavior such as eye gaze, body movement, gestures, and proxemics offer key insights into the dynamics of social interactions. Eye gaze, for example, is related to action patterns (Kendon and Cook, 1969) and attention (Kütt et al., 2020; Semsar et al., 2020), and provides important signals to understanding dyadic collaborations (Andrist et al., 2015; Jing et al., 2022), user intentions and coordination (Amati and Brennan, 2018), learning outcomes (Pi et al., 2020), conversational attention (Vertegaal et al., 2001), and social engagement (Nakano and Ishii, 2010). People’s gestures and body orientation can influence the sense of feeling addressed by conversation partners (Nagels et al., 2015), while their proxemics and body orientation, in addition to eye gaze, can be effective in predicting one’s intention of joining social groups (Bönsch et al., 2020).
@@
It is then unsurprising that nonverbal communication makes up a key component of digital social interactions, for which social VR offers additional benefits compared to video conferencing. By tracking the users’ eyes, heads, and hands, VR systems render avatars in virtual spaces that more accurately convey nonverbal cues such as interpersonal distance and eye gaze, fostering social interactions that are perceived positively and comparable to face-to-face interactions (Maloney et al., 2020; Smith and Neff, 2018). Communication through nonverbal behaviors such as bodily and facial gestures induce more positive dyadic interactions in VR, with these cues also being predictive of interpersonal attraction (Oh Kruzic et al., 2020). Using motion data, VR social interactions can also be augmented, for example through increasing mimicry (Tarr et al., 2018; Raffard et al., 2018; Bailenson and Yee, 2005) and transforming eye gaze and interpersonal distance (Roth et al., 2018, 2019; Wang et al., 2024b, 2023). With VR placing users in controllable virtual environments irrespective of their physical context, it is also possible to examine nonverbal behavior longitudinally (Miller et al., 2023a, b). These works highlighted nonverbal behavior changes over time – participants in social VR looked at others more and interpersonal distance increased over weeks (Miller et al., 2023a), and the identifiability of users lowered when there is greater temporal delay between training and testing sessions (Miller et al., 2023b).
@@
VR tracking data also enables fine-grain analyses on topics such as synchrony (Sun et al., 2019; Miller et al., 2021), self-efficacy and learning (Queiroz et al., 2023), physiological responses (Luong and Holz, 2022), classroom discourse (Stark et al., 2024), interpersonal distance (Miller et al., 2023a; DeVeaux et al., ), context (Han et al., 2024), design behaviors (Wang et al., 2024a), and user identification (Miller et al., 2020; Nair et al., 2023; Miller et al., 2022, 2023b; Moore et al., 2021). Particularly relevant to our work is the research of DeVeaux et al. (DeVeaux et al., ), which leveraged nonverbal and verbal tracking data to study language use. The authors extracted linguistic patterns from verbal transcripts and notably found a positive correlation between the use of impersonal pronouns and the median interpersonal distance. This research highlighted the importance of considering language use to study VR affordances and the potential of leveraging nonverbal behavior to uncover additional psychological nuances surveys alone may fail to capture. Key distinctions between their work and ours are their focus on the nonverbal behavior of interpersonal distance, linguistic styles, and session-level analyses. In contrast, we studied turn-taking behavior characteristics (e.g., who will speak next, when they will speak) and focused primarily on predicting behaviors at moments within sessions. Beyond interpersonal distance, we also included other nonverbal features (e.g., gaze, egocentric motion) to capture social dynamics. One other distinction lies in the VR activities analyzed: while DeVeaux et al. (DeVeaux et al., ) studied instructor-led student group discussions collected during a university course in Fall 2021, we focused on open-ended VR group activities that were recorded a year later through the same course.
@@
Concretely, we contribute to past works by leveraging motion data collected in VR social settings to predict turn-taking behaviors. Findings on changes in nonverbal behaviors over time (Miller et al., 2023a, b) have emphasized the need to study social behaviors through a longitudinal lens and across multiple sessions. These insights motivated our examination of the robustness and generalizability of turn-taking predictions across unseen weeks, groups, and activities.
@@ 2.2. Modeling Human Behavior and Individual- and Group- Level Differences using Tracking Data
Besides using tracking data to understand human behaviors, monitoring motion and user data in real-time, often in an unobtrusive manner, can further enable synchronous interventions and assessments. For example, as smartphone data such as touch and typing behaviors (Wampfler et al., 2022; Kovačević et al., 2023) and audio, textual, and video data (Kampman et al., 2018) are predictive of personality traits and affective states, one can construct more personalized experiences based on the user’s current behaviors. Another example involves monitoring eye gaze to assess cognitive load, and using this information to facilitate real-time adaptations of mixed reality interfaces (Lindlbauer et al., 2019). More recently, researchers have also built models for predicting team viability using online team text conversations and demonstrated the potential of using automated features for real-time assessments and intervention (Cao et al., 2021). Others have proposed using individual and social behaviors to model purchase decisions for improving e-commerce recommendations (Xu et al., 2019) and job burnouts for informing early-stage interventions (Wu et al., 2021a).
@@
Similarly motivated, we see opportunities in predicting VR speech behaviors for real-time interventions and assistance (e.g., instantiating social behaviors for virtual assistants, reducing unwanted overlapping speech). Many have explored how tracking data can be used to predict turn-taking behaviors such as the next speaker in a group setting. Fortunately, though human subjects struggled to infer speaker behavior from audio alone (Heeman and Lunsford, 2015), research that leverages tracking data for speech behavior predictions has shown greater promise. Works looked to speech sequences (Parker, 1988), mouth motion (Ishii et al., 2016a), head movement (Ishii et al., 2017), respiration (Ishii et al., 2016c), gaze behavior (Ishii et al., 2016b; Lee et al., 2023), and more broadly user motion (Chen et al., 2024) as predictors when modeling speech behaviors. Notably, these works have found that markov models, a type of stochastic model that predicts future states based on past ones, were effective in predicting the next speaker based on the two preceding speakers (Parker, 1988) and that support vector machines leveraging features on gaze transition patterns are predictive of turn-changing characteristics (Ishii et al., 2016b).
@@
Unlike the social scenes examined previously, where social groups completed a single or series of similar tasks (Chen et al., 2024; Ishii et al., 2016b) and either stood in fixed positions (Lee et al., 2023) or remained seated (Chen et al., 2024; Ishii et al., 2016b), VR social interactions are diverse in activities and typically involve less constraints on user movement in both physical and virtual spaces. These differences raise the question of whether turn-taking behavior predictions are possible when there are greater variations in activities and nonverbal cues such as interpersonal distance and eye gaze. Failure to make reliable and robust predictions in these scenarios limits the prospects of real-time intervention and assistance in immersive settings. To address this, we investigated VR turn-taking behaviors of social groups engaging in open-ended group activities across four weeks with no restriction on virtual user motion. We contribute to past literature by also interpreting how features are related to turn-taking predictions, highlighting their theoretical implications and practical significance for future applications.

### v2

@@ Abstract.
In networked virtual reality (VR), user behaviors, individual differences, and group dynamics can serve as important signals into future speech behaviors, such as who the next speaker will be and the timing of turn-taking behaviors. The ability to predict and understand these behaviors offers opportunities to provide adaptive and personalized assistance, for example helping users with varying sensory abilities navigate complex social scenes and instantiating virtual moderators with natural behaviors. In this work, we predict turn-taking behaviors using features extracted based on social dynamics literature. We discuss results from a large-scale VR classroom dataset consisting of 77 sessions and 1660 minutes of small-group social interactions collected over four weeks. In our evaluation, gradient boosting classifiers achieved the best performance, with accuracies of 0.71–0.78 AUC (area under the ROC curve) across three tasks concerning the “what”, “who”, and “when” of turn-taking behaviors. In interpreting these models, we found that group size, listener personality, speech-related behavior (e.g., time elapsed since the listener’s last speech event), group visual attention (e.g., the group’s head orientation towards the speaker), and the listener and previous speaker’s head pitch, head y-axis position, and left hand y-axis position more saliently influenced predictions. Results suggested that these features remain reliable indicators in novel social VR settings, as prediction performance is robust over time and with groups and activities not used in the training dataset. We discuss theoretical and practical implications of the work.
@@ 1. Introduction
Effective and organized social interactions require its social actors to accurately infer speaking intention and take turns accordingly. In most platforms for facilitating computer mediated-communication (CMC), users retain a high level of autonomy over when and how they speak, utilizing features such as the muting and unmuting button during videoconferencing. However, along with the benefits of user control comes the responsibility for individuals to correctly infer social group dynamics. Such inferences can be difficult to navigate as users utilize digital platforms such as video conferencing and social virtual reality (VR), where the level of immersion and representation of users (e.g., nonverbal behavior, avatar appearance) can differ from one another (Steinicke et al., 2020; Abdullah et al., 2021). In immersive social interactions, differences in avatar representation and characteristics of social VR platforms can yield adverse outcomes such as unwanted overlap in speech (Barreda-Ángeles et al., 2023; Williamson et al., 2021), lack of control for users with physical disabilities (Maloney et al., 2020), misperceived and absence of nonverbal cues (Maloney et al., 2020; Moustafa and Steed, 2018), and difficulty gauging the end of speaking turns (Moustafa and Steed, 2018; Williamson et al., 2021). For users with visual and audio disabilities, it can also be challenging for them to engage in seamless interactions based solely off of perceptual and auditory feedback (Kushalnagar and Vogler, 2020; Anderson, 2021). Paradoxically, as immersive communication evolves to be hardware-agnostic, with users interacting across traditional screen devices and mixed-reality headsets (Qiu et al., 2023), the mismatch in social cues delivered and perceived by users across different devices present further complications for users to navigate socially.
Being able to predict user intentions for actions such as turn taking can alleviate these challenges through real-time interventions and training. Social VR platforms can implement personalized assistance for users who are struggling to interpret nonverbal cues under technological constraints (e.g., joining large meetings though audio feed) or are different in physical abilities (e.g., auditory and visual abilities). This assistance can also benefit populations with medical conditions such as autism and ADHD that inhibit their ability to accurately interpret nonverbal behaviors (Freeman et al., 2002; Hall et al., 1999). Relatedly, understanding what behavior precedes turn taking can be useful in training practitioners such as educators and healthcare workers to better navigating social settings. For example, by training early-career practitioners to recognize nonverbal cues and individually characteristics influencing speaking intentions, they can be better equipped to navigate social situations.
@@
In more structured settings such as classrooms and focus groups, knowing who will likely speak next and what nonverbal cues precede speech can improve social dynamics. For one, practitioners who need to monitor multiple virtual sessions simultaneously, inevitably missing nonverbal and verbal cues, can leverage turn-taking predictions to remain informed of the social dynamics and better facilitate group conversations. These practitioners can also instantiate virtual agents to moderate social VR scenes with natural nonverbal behaviors that do not cut off current speakers and more broadly design mechanisms for smoothly guiding the conversations. Failure to accurately infer social dynamics can yield unwanted interruptions and lower the perceived ease of use of these intelligent systems (Maier et al., 2022). As such, we argue that it is important to study whether we can robustly predict virtual turn-taking behaviors and how nonverbal and verbal behaviors as well as individual and group characteristics influence turn-taking predictions.
@@
In this work, we investigated turn-taking behaviors in VR and leveraged the medium’s fine-grain tracking data on motion and verbal behaviors. Using the VR motion data, researchers have implemented pipelines for deriving features such as head movement (Miller et al., 2023a) and gaze (Abdullah et al., 2021), both of which have been previously useful for predicting speech behaviors in face-to-face social interactions (Ishii et al., 2017, 2016b). As such, our first research question considers whether it is possible to predict turn taking.
@@
Research Question 1 (RQ1): Can we predict turn-taking behaviors in VR open-ended group activities from features extracted from motion tracking data, speech-related behavior, and individual and group differences?
Building on this, we leveraged VR to place individuals into varying virtual spaces and studied social dynamics across groups, social activities, and spatial contexts over time. Studying diverse social interactions at scale is important as building models that are robust to unseen groups, activities, and time suggests higher practicality and reliability. Concretely, our second research question concerns generalizability.
@@
Finally, we interpreted the predictive models. Analyzing how features predict turn taking can deepen our understanding of virtual social dynamics, allowing future researchers to benchmark against face-to-face interactions. This understanding can also help practitioners build more efficient models with curated features and compose social groups exhibiting certain turn-taking behaviors.
Research Question 3 (RQ3): How are the extracted features related to turn-taking predictions and performance? Specifically, what features are strongly related to performance, and how are they associated with turn-taking predictions?
To answer these questions, we built models predicting speech dynamics, focusing on three tasks that capture the “what”, “who”, and “when” of turn-taking behaviors. We formulated features based on tracking data and individual and group characteristics based on social dynamics literature. We evaluated performance across 77 sessions and 1660 minutes of open-ended VR group discussions spanning four weeks, each with three to four university students. Notably, we found that gradient boosting classifiers achieved the best accuracies, predicting turn transitions and identifying the new speaker with 0.75–0.78 AUC (i.e., area under the ROC curve), and differentiating between moments immediately preceding turn transitions and those sampled at prior moments with accuracies of 0.71–0.72 AUC. These results demonstrated that we were able to predict turn-taking behaviors with accuracies considerably higher than prediction by chance (i.e., 0.50 AUC) and highlighted the potential non-linear relationships and interactions between extracted features.
Additional feature analyses revealed the importance of listener personality, group size, speech-related behavior (e.g., preceding speaker sequence), group visual attention, head pitch and y-axis position, and left hand y-axis position. Our models showed comparable results when evaluated on time, activities, and groups not seen during training, which demonstrated the reliability of these turn-taking indicators in novel social settings. From these results, we highlight their theoretical relevance and outline practical implications for how practitioners can leverage behavioral predictions to facilitate effective VR social interactions.
In summary, we make the following contributions. First, we formulated features based on social dynamics literature and leveraged both the VR tracking data and individual and group characteristics. Using a large-scale VR dataset with 1660 minutes of open-ended activities collected over four weeks, we then demonstrated the feasibility and robustness of predicting VR turn-taking behaviors. Finally, through interpreting how features influence model performance and prediction probabilities, we present theoretical and practical insights and highlight the potential of our work for intervention, support, and training.
@@ 2.1. Nonverbal Behavior in Social Interactions
Nonverbal behavior offers insights into social dynamics. Gaze, for example, is related to action patterns (Kendon and Cook, 1969) and attention (Kütt et al., 2020; Semsar et al., 2020), and provides important signals to understanding dyadic collaborations (Andrist et al., 2015; Jing et al., 2022), user intention and coordination (Amati and Brennan, 2018), learning outcomes (Pi et al., 2020), conversational attention (Vertegaal et al., 2001), and social engagement (Nakano and Ishii, 2010). Gesture and body orientation can influence the sense of feeling addressed by conversation partners (Nagels et al., 2015), while proxemics, body orientation, and gaze, can be effective in predicting one’s intention of joining social groups (Bönsch et al., 2020).
@@
It is then unsurprising that nonverbal communication makes up a key component of virtual interactions, for which social VR offers additional benefits compared to video conferencing. By tracking the users’ eyes, heads, and hands, VR systems render avatars that accurately convey nonverbal cues such as interpersonal distance and gaze, fostering social interactions that are perceived positively and comparable to face-to-face interactions (Maloney et al., 2020; Smith and Neff, 2018). Communication through nonverbal behaviors such as bodily and facial gestures induce more positive dyadic interactions in VR, with these cues also being predictive of interpersonal attraction (Oh Kruzic et al., 2020). Using motion data, VR social interactions can also be augmented, for example through increasing mimicry (Tarr et al., 2018; Raffard et al., 2018; Bailenson and Yee, 2005) and transforming gaze and interpersonal distance (Roth et al., 2018, 2019; Wang et al., 2024b, 2023). With VR placing users in controllable virtual environments, it is also possible to examine nonverbal behavior longitudinally (Miller et al., 2023a, b). Works highlighted how nonverbal behavior changes over time – users looked at others more and their interpersonal distance increased (Miller et al., 2023a), and user identifiability lowered with greater temporal delay between training and testing sessions (Miller et al., 2023b).
@@
VR tracking data also enables fine-grain analyses on synchrony (Sun et al., 2019; Miller et al., 2021), self-efficacy and learning (Queiroz et al., 2023), physiological responses (Luong and Holz, 2022), classroom discourse (Stark et al., 2024), interpersonal distance (Miller et al., 2023a; DeVeaux et al., ), context (Han et al., 2024), design behaviors (Wang et al., 2024a), and user identification (Miller et al., 2020; Nair et al., 2023; Miller et al., 2022, 2023b; Moore et al., 2021). Particularly relevant to our work is the research of DeVeaux et al. (DeVeaux et al., ). The authors extracted linguistic patterns from transcripts and notably found a positive correlation between the use of impersonal pronouns and the median interpersonal distance. This research highlighted the importance of considering language use to study VR affordances and the potential of leveraging nonverbal behavior to uncover psychological nuances that surveys fail to capture. Key distinctions between their work and ours are their focus on the nonverbal behavior of interpersonal distance, linguistic styles, and session-level analyses. In contrast, we studied turn-taking behaviors and focused on predicting them at moments within sessions. Beyond interpersonal distance, we also included other nonverbal features (e.g., egocentric motion). One other distinction lies in the VR activities analyzed: while DeVeaux et al. (DeVeaux et al., ) studied instructor-led discussions collected during a university course in Fall 2021, we focused on topen-ended activities recorded a year later through the same course.
@@
We contribute to past works by predict turn-taking behaviors through VR motion data. Findings on changes in nonverbal behaviors over time (Miller et al., 2023a, b) emphasized the need to study social behaviors through a longitudinal lens and across multiple sessions. These insights motivated our examination of turn-taking predictions across unseen weeks, groups, and activities.
@@ 2.2. Modeling Human Behavior and Individual- and Group- Level Differences using Tracking Data
Besides using tracking data to understand human behaviors, monitoring user motion, often unobtrusively, can further enable interventions and assessments. For example, as smartphone data such as touch and typing behaviors (Wampfler et al., 2022; Kovačević et al., 2023) and audio, textual, and video data (Kampman et al., 2018) are predictive of personality traits and affective states, one can construct personalized experiences based on the user’s current behaviors. Another example involves monitoring gaze to assess cognitive load, and using this information to facilitate adaptations of mixed reality interfaces (Lindlbauer et al., 2019). Recently, researchers also built models for predicting team viability using online team text conversations and demonstrated the potential of using automated features for assessments and intervention (Cao et al., 2021). Others proposed using individual and social behaviors to model purchase decisions for e-commerce recommendations (Xu et al., 2019) and job burnouts for early-stage interventions (Wu et al., 2021a).
@@
Similarly, we see opportunities in predicting VR speech behaviors for intervention and assistance. Fortunately, though humans can struggle to infer speaker behavior (Heeman and Lunsford, 2015), research that leverages tracking data for speech behavior predictions has shown greater promise. Works looked to speech sequences (Parker, 1988), mouth and head motion (Ishii et al., 2016a, 2017), respiration (Ishii et al., 2016c), gaze behavior (Ishii et al., 2016b; Lee et al., 2023), and more broadly user motion (Chen et al., 2024) as predictors when modeling speech behavior. Notably, these works have found that markov models were effective in predicting the next speaker based on the two preceding speakers (Parker, 1988) and that support vector machines leveraging features on gaze transition patterns are predictive of turn-changing characteristics (Ishii et al., 2016b).
@@
Unlike the social scenes examined in previous studies, where users completed a single or series of similar tasks (Chen et al., 2024; Ishii et al., 2016b) and either stood in fixed positions (Lee et al., 2023) or remained seated (Chen et al., 2024; Ishii et al., 2016b), VR social interactions are diverse in activities and typically involve less constraints on movement in the physical and virtual spaces. This raises the question of whether turn-taking predictions are possible when there are greater variations in activities and nonverbal cues. Failure to make robust predictions in these scenarios limits the prospects of intervention and assistance in VR. To address this, we investigated VR turn-taking behaviors from four weeks of open-ended activities with little restriction on virtual motion. We contribute by also interpreting how features are related to turn-taking predictions, highlighting their theoretical and practical significance.